# A Hausdorff-measure boundary element method for acoustic scattering by fractal screens

### v1

@@ Appendix A Besov spaces on $d$-sets
In this appendix we show that the spaces denoted by $B_{\alpha}^{p,q}(\Gamma)$, and defined in terms of atoms in [30, section 6], coincide with the spaces denoted by $\mathbb{B}_{p,q,0}^{\alpha}(\Gamma)$ defined in [8, Def. 6.3], under the assumptions $\alpha>0$, $1\leq p,q<\infty$, and $\Gamma$ being a $d$-set, $0<d<n$, preserving Markov’s inequality in the sense of [30, §4] (or see [31, p. 34], and note Remark A.6 below). As a consequence (see Corollary A.4) we have that our space $\mathbb{H}^{\alpha}(\Gamma)$, defined in §2.4, coincides with $B_{\alpha}^{2,2}(\Gamma)$ under the same assumptions on $\alpha$ and $\Gamma$. [^19] We recall (see the discussion in §2.4 and [8, Rem. 6.4]) that, for $0<\alpha<1$, $\mathbb{H}^{\alpha}(\Gamma)$ also coincides with the Besov space $B^{\alpha}_{2,2}(\Gamma)$ of [31].
@@
The following definition of $B_{\alpha}^{p,q}(\Gamma)$ is the second, equivalent, definition of [30, §6].
@@
Let $\alpha>0$, $1\leq p,q<\infty$, and $\Gamma$ be a $d$-set, $0<d<n$, preserving Markov’s inequality. Then the space $\mathbb{B}_{p,q,0}^{\alpha}(\Gamma)$ defined in [8, Def. 6.3] coincides with the space $B_{\alpha}^{p,q}(\Gamma)$ of Definition A.1 above, with equivalence of norms.
@@
On the other hand, since $K>\alpha+\frac{n-d}{p}$ and $0>-\alpha-\frac{n-d}{p}$, it follows from [7, Thm. 2.3, Rem. 2.4 and Thm. 2.5] and part 1 of the above Remark A.2 that
@@
for some constant $c>0$ independent of $f$. From elementary embeddings and the linearity and continuity of the trace operator $\mathrm{tr}_{\Gamma}$ from [8, Prop. 6.2],
@@
the latter equality being justified by the arguments in [8, Rem. 6.4] and the fact that each $a^{\nu m}$ is continuous. But then $\mathrm{tr}_{\Gamma}g$ must be the same as $f$ (recall the representation (130) and the fact that $a^{\nu m}|_{\Gamma}$ and $a^{\nu m}$ define the same class in $\mathbb{L}_{p}(\Gamma)$. So, we have proved that $f$ also belongs to $\mathbb{B}_{p,q,0}^{\alpha}(\Gamma)$. Additionally, using (132),
@@
Since $K+1>\alpha+\frac{n-d}{p}$ and $0>-\alpha-\frac{n-d}{p}$, it follows from [7, Thm. 2.3 and Thm. 2.5] that
@@
for some constant $c^{\prime}>0$ independent of $f$. Again using elementary embeddings and the linearity and continuity of the trace operator $\mathrm{tr}_{\Gamma}$ from [8, Prop. 6.2],
@@
Our space $\mathbb{H}^{\alpha}(\Gamma)$ coincides with the spaces $\mathbb{H}^{\alpha}_{2,0}(\Gamma)=\mathbb{F}^{\alpha}_{2,2,0}(\Gamma)$ defined in [8, Def. 6.3]. Moreover, since (for all $s\in\mathbb{R}$) the Triebel-Lizorkin space $F_{2,2}^{s}{(\mathbb{R}^{n})}$ coincides with the Besov space $B_{2,2}^{s}{(\mathbb{R}^{n})}$, with equivalent norms, it follows, inspecting the definitions of $\mathbb{F}^{\alpha}_{2,2,0}(\Gamma)$ and $\mathbb{B}_{2,2,0}^{\alpha}(\Gamma)$ in [8, Def. 6.3], that $\mathbb{F}^{\alpha}_{2,2,0}(\Gamma)=\mathbb{B}_{2,2,0}^{\alpha}(\Gamma)$, also with equivalence of norms, and the result follows immediately from the previous proposition. ∎
@@
Characterizations of trace spaces in terms of atomic representations, like the one given in Proposition A.3 above, are not new. With a somewhat different approach and in the larger setting of the so-called $h$-sets, they can already be seen in [6, Prop. 3.5.4].
@@
If $\Gamma\subset\mathbb{R}^{n}$ is a $d$-set and d>n1d>n-1, then the hypothesis that $\Gamma$ preserves Markov’s inequality is automatically satisfied [31, Thm. 3 on p. 39].
@@ Appendix B Inverse estimates
In this appendix we prove that the inverse estimate (86), established for $0<t<1$ in Theorem 5.10, holds in fact for the extended range $0<t\leq J$, for any $J\in\mathbb{N}$, moreover with a constant in the inverse estimate that is independent of $t$. Our proof is modelled on standard inverse estimate arguments for negative exponent Sobolev spaces, e.g. [20, Theorem 4.6], with the important difference that it is, of course, impossible to support the smooth “bubble functions” (in the terminology of [20, §4.3]), that are a key tool in the arguments of [20] and in the proof of earlier inverse estimate results, inside an element $\Gamma_{{\boldsymbol{m}}}\subset\Gamma$, as $\Gamma$ has empty interior. In the case that $\Gamma$ is a disjoint IFS attractor, it turns out that, to carry through an analogous argument, it is enough to replace the bubble function by a smooth function supported in a carefully chosen neighbourhood of $\Gamma_{{\boldsymbol{m}}}$ (the function $\sigma_{{\boldsymbol{m}}}\in C_{0}^{\infty}(\mathbb{R}^{n})$ in the proof below). The assumption that $\Gamma$ is a disjoint IFS attractor implies that $\Gamma$ is a $d$-set with $0<d<n$ (see Lemma 2.6), but, in contrast to Theorem 5.10, we make no additional constraint on $d$. We continue to assume in this appendix that $h$ lies in the range (78), i.e. that
@@
Since $\Gamma_{1}$, …, $\Gamma_{M}$ are disjoint, $\Gamma$ satisfies the OSC for some open set $O\supset\Gamma$, by Lemma 2.5. By a standard mollification we can construct a $\sigma\in C_{0}^{\infty}(\mathbb{R}^{n})$ supported in $O$ such that $\sigma=1$ in a neighbourhood of $\Gamma$. For each ${\boldsymbol{m}}=(m_{1},...,m_{\ell})\in L_{h}$, let $\rho_{{\boldsymbol{m}}}:=\prod_{j=1}^{\ell}\rho_{m_{j}}$,
@@
For $j\in\mathbb{N}_{0}$ and ${\boldsymbol{m}}\in L_{h}$, using standard properties of Fourier transforms (e.g., [26, Prop. 2.2.11]),
@@
Acknowledgements. We are grateful to Martin Costabel for instructive discussions related to this project. SC-W was supported by Engineering and Physical Sciences Research Council (EPSRC) Grant EP/V007866/1, and DH and AG by EPSRC grants EP/S01375X/1 and EP/V053868/1. AM was supported by the PRIN project “NA-FROM-PDEs” and by the MIUR through the “Dipartimenti di Excellenza” Program (2018–2022)–Dept. of Mathematics, University of Pavia. AC was supported by CIDMA (Center for Research and Development in Mathematics and Applications) and FCT (Foundation for Science and Technology) within project UIDB/04106/2020.

### v2

@@ 6.5 Non-homogeneous or non-disjoint IFS attractors
7 Conclusions and Future Work
In this paper we presented and analysed a piecewise-constant Galerkin BEM for acoustic scattering in $\mathbb{R}^{n+1}$, $n=1,2$, by a sound-soft planar screen $\Gamma\subset\Gamma_{\infty}=\mathbb{R}^{n}\times\{0\}$. Our BEM is defined whenever the screen $\Gamma$ is a compact $d$-set, for some $n-1<d\leq n$, which includes cases where $\Gamma$ is fractal. It is based on an integral equation formulation in which integration is carried out with respect to Hausdorff measure $\mathcal{H}^{d}$. For any compact $d$-set we proved that the method converges as the mesh width $h$ tends to zero (see Theorem 5.1). Regarding the relationship between mesh width and wavelength, in our numerical results presented in §6 we observe the same behaviour one obtains for a conventional BEM on a smooth scatterer: as $h$ decreases towards zero there is a pre-asymptotic phase until $h$ reaches the wavelength scale, beyond which one observes the predicted asymptotic behaviour. In the case where $\Gamma$ is the disjoint attractor of an IFS satisfying the OSC, the “elements” in the BEM are self-similar subsets of $\Gamma$, and in §5.4 we showed how the Galerkin integrals can be evaluated using quadrature rules from [30]. In this case we also proved fully-discrete convergence rates, under certain regularity assumptions on the integral equation solution (see Theorem 5.6, Corollaries 5.17 and 5.18). Specifically, we showed that the BEM solution converges like $O(h^{s+1/2})$ as $h\to 0$, and the near- and far-field solutions like $O(h^{2s+1})$, assuming $\phi\in H^{s}_{\Gamma}$ for some s>1/2s>-1/2. We proved the existence of such an $s$ in Remark 5.7, using Proposition 4.9, but our numerical results in §6 suggest that, for sufficiently smooth data, it may hold that $\phi\in H^{s}_{\Gamma}$ for all s<(nd)/2s<-(n-d)/2, i.e. for all $s$ such that the space $H^{s}_{\Gamma}$ is non-trivial. Guided by these observations, we formulated Conjecture 4.8, which is a statement about the range of Sobolev spaces on which the Hausdorff-measure integral operator $\mathbb{S}$ (defined in (15)) is invertible.
Proving or disproving Conjecture 4.8 is the main outstanding theoretical question relating to the paper. Other avenues for future research include:
  • The extension of our analysis to non-disjoint fractal screens, such as the Sierpinski triangle screen considered in §6.5. Singular quadrature rules for such cases (generalising those described in §5.4) have been presented recently in [29], but extending our BEM convergence analysis will require a suitable generalisation of the wavelet approximation theory of [36], which is yet to be worked out.
  • The extension of our convergence rate analysis to the case $d=n$. This requires somewhat different techniques to the case $d<n$, and will be presented in a separate article [11].
  • The generalisation of our Hausdorff BEM to scattering by non-planar fractal structures. Results in this direction were presented recently in [10].
  • The extension to Neumann problems. Theoretical results relating to integral equation formulations of Neumann screen problems were presented in [15]. These show that the Neumann boundary condition is “weaker” than the Dirichlet condition, in the sense that Neumann screens do not scatter waves unless $H^{1/2}_{\Gamma}\neq\{0\}$, which requires in particular that $\Gamma\subset\Gamma_{\infty}\cong\mathbb{R}^{n}$ has positive $n$-dimensional Lebesgue measure. Hence, in the context of scattering by screens that are attractors of IFS satisfying the OSC, only the case $d=n$ is relevant. A specific example would be scattering by a sound-hard Koch snowflake screen. However, the development of a BEM for such Neumann problems using a mesh of fractal elements (as we consider in the current paper for the Dirichlet case) is complicated by the fact that non-trivial piecewise polynomials on such meshes cannot be continuous, and hence cannot be $H^{1/2}_{\Gamma}$-conforming. Thus a different approach is required, perhaps involving a discontinuous Galerkin discretization. We remark that a rigorous convergence analysis for a prefractal BEM approach to the related impedance problem was presented in [3].
  • A more detailed investigation into the relative accuracy of our Hausdorff BEM and the alternative prefractal BEM of [19], extending the preliminary analysis of §6.4. For many problems the Hausdorff BEM appears to be more accurate than the prefractal BEM. However, the comparison between the two is rather subtle: for the examples we considered in §6.4, while both methods appear to converge at the same rate, for the same number of degrees of freedom the Hausdorff BEM can be over 1000 times more accurate than the prefractal BEM, or 100 times less accurate, depending on the fractal simulated. As yet we do not understand why.
  • The combination of our Hausdorff BEM with accelerated linear algebra via a technique such as $\mathcal{H}$-matrix compression or the fast multipole method. This would allow the simulation of larger problems with finer mesh widths, permitting the calculation of higher accuracy solutions and/or the study of higher frequency problems.
@@ Appendix A Besov spaces on $d$-sets
In this appendix we show that the spaces denoted by $B_{\alpha}^{p,q}(\Gamma)$, and defined in terms of atoms in [36, section 6], coincide with the spaces denoted by $\mathbb{B}_{p,q,0}^{\alpha}(\Gamma)$ defined in [9, Def. 6.3], under the assumptions $\alpha>0$, $1\leq p,q<\infty$, and $\Gamma$ being a $d$-set, $0<d<n$, preserving Markov’s inequality in the sense of [36, §4] (or see [37, p. 34], and note Remark A.6 below). As a consequence (see Corollary A.4) we have that our space $\mathbb{H}^{\alpha}(\Gamma)$, defined in §2.4, coincides with $B_{\alpha}^{2,2}(\Gamma)$ under the same assumptions on $\alpha$ and $\Gamma$. [^18] We recall (see the discussion in §2.4 and [9, Rem. 6.4]) that, for $0<\alpha<1$, $\mathbb{H}^{\alpha}(\Gamma)$ also coincides with the Besov space $B^{\alpha}_{2,2}(\Gamma)$ of [37].
@@
The following definition of $B_{\alpha}^{p,q}(\Gamma)$ is the second, equivalent, definition of [36, §6].
@@
Let $\alpha>0$, $1\leq p,q<\infty$, and $\Gamma$ be a $d$-set, $0<d<n$, preserving Markov’s inequality. Then the space $\mathbb{B}_{p,q,0}^{\alpha}(\Gamma)$ defined in [9, Def. 6.3] coincides with the space $B_{\alpha}^{p,q}(\Gamma)$ of Definition A.1 above, with equivalence of norms.
@@
On the other hand, since $K>\alpha+\frac{n-d}{p}$ and $0>-\alpha-\frac{n-d}{p}$, it follows from [8, Thm. 2.3, Rem. 2.4 and Thm. 2.5] and part 1 of the above Remark A.2 that
@@
for some constant $c>0$ independent of $f$. From elementary embeddings and the linearity and continuity of the trace operator $\mathrm{tr}_{\Gamma}$ from [9, Prop. 6.2],
@@
the latter equality being justified by the arguments in [9, Rem. 6.4] and the fact that each $a^{\nu m}$ is continuous. But then $\mathrm{tr}_{\Gamma}g$ must be the same as $f$ (recall the representation (128) and the fact that $a^{\nu m}|_{\Gamma}$ and $a^{\nu m}$ define the same class in $\mathbb{L}_{p}(\Gamma)$. So, we have proved that $f$ also belongs to $\mathbb{B}_{p,q,0}^{\alpha}(\Gamma)$. Additionally, using (130),
@@
Since $K+1>\alpha+\frac{n-d}{p}$ and $0>-\alpha-\frac{n-d}{p}$, it follows from [8, Thm. 2.3 and Thm. 2.5] that
@@
for some constant $c^{\prime}>0$ independent of $f$. Again using elementary embeddings and the linearity and continuity of the trace operator $\mathrm{tr}_{\Gamma}$ from [9, Prop. 6.2],
@@
Our space $\mathbb{H}^{\alpha}(\Gamma)$ coincides with the spaces $\mathbb{H}^{\alpha}_{2,0}(\Gamma)=\mathbb{F}^{\alpha}_{2,2,0}(\Gamma)$ defined in [9, Def. 6.3]. Moreover, since (for all $s\in\mathbb{R}$) the Triebel-Lizorkin space $F_{2,2}^{s}{(\mathbb{R}^{n})}$ coincides with the Besov space $B_{2,2}^{s}{(\mathbb{R}^{n})}$, with equivalent norms, it follows, inspecting the definitions of $\mathbb{F}^{\alpha}_{2,2,0}(\Gamma)$ and $\mathbb{B}_{2,2,0}^{\alpha}(\Gamma)$ in [9, Def. 6.3], that $\mathbb{F}^{\alpha}_{2,2,0}(\Gamma)=\mathbb{B}_{2,2,0}^{\alpha}(\Gamma)$, also with equivalence of norms, and the result follows immediately from the previous proposition. ∎
@@
Characterizations of trace spaces in terms of atomic representations, like the one given in Proposition A.3 above, are not new. With a somewhat different approach and in the larger setting of the so-called $h$-sets, they can already be seen in [7, Prop. 3.5.4].
@@
If $\Gamma\subset\mathbb{R}^{n}$ is a $d$-set and d>n1d>n-1, then the hypothesis that $\Gamma$ preserves Markov’s inequality is automatically satisfied [37, Thm. 3 on p. 39].
@@ Appendix B Inverse estimates
In this appendix we prove that the inverse estimate (84), established for $0<t<1$ in Theorem 5.10, holds in fact for the extended range $0<t\leq J$, for any $J\in\mathbb{N}$, moreover with a constant in the inverse estimate that is independent of $t$. Our proof is modelled on standard inverse estimate arguments for negative exponent Sobolev spaces, e.g. [24, Theorem 4.6], with the important difference that it is, of course, impossible to support the smooth “bubble functions” (in the terminology of [24, §4.3]), that are a key tool in the arguments of [24] and in the proof of earlier inverse estimate results, inside an element $\Gamma_{{\boldsymbol{m}}}\subset\Gamma$, as $\Gamma$ has empty interior. In the case that $\Gamma$ is a disjoint IFS attractor, it turns out that, to carry through an analogous argument, it is enough to replace the bubble function by a smooth function supported in a carefully chosen neighbourhood of $\Gamma_{{\boldsymbol{m}}}$ (the function $\sigma_{{\boldsymbol{m}}}\in C_{0}^{\infty}(\mathbb{R}^{n})$ in the proof below). The assumption that $\Gamma$ is a disjoint IFS attractor implies that $\Gamma$ is a $d$-set with $0<d<n$ (see Lemma 2.6), but, in contrast to Theorem 5.10, we make no additional constraint on $d$. We continue to assume in this appendix that $h$ lies in the range (76), i.e. that
@@
Since $\Gamma_{1}$, …, $\Gamma_{M}$ are disjoint, $\Gamma$ satisfies the OSC for some open set $O\supset\Gamma$ by Lemma 2.5. By a standard mollification we can construct a $\sigma\in C_{0}^{\infty}(\mathbb{R}^{n})$ supported in $O$ such that $\sigma=1$ in a neighbourhood of $\Gamma$. For each ${\boldsymbol{m}}=(m_{1},...,m_{\ell})\in L_{h}$, let $\rho_{{\boldsymbol{m}}}:=\prod_{j=1}^{\ell}\rho_{m_{j}}$,
@@
For $j\in\mathbb{N}_{0}$ and ${\boldsymbol{m}}\in L_{h}$, using standard properties of Fourier transforms (e.g., [31, Prop. 2.2.11]),
@@
Appendix C Table of definitions
Symbol/terminology Defined/introduced
$\Gamma$, $\Gamma_{\infty}$ §1
$\mathcal{H}^{\alpha}$, $\rm dim_{H}$}, $|E|$ (Lebesgue measure), $d$-set §2.1
IFS, OSC, homogeneous, self-similar, disjoint §2.3
[…]
$\mathbb{L}_{2}(\Gamma)$, $\mathbb{L}_{\infty}(\Gamma)$, $\mathbb{H}^{t}(\Gamma)$, $\mathrm{tr}_{\Gamma}$, $\mathrm{tr}_{\Gamma}^{*}$ §2.4 & §2.5
$W^{1}(\Omega)$, $W^{1,{\rm loc}}(\Omega)$, $\gamma^{\pm}$, $\partial_{\boldsymbol{n}}^{\pm}$, $C^{\infty}_{0,\Gamma}$ §2.5
$_{\ell}, II_{\mathbb{N}}$, $I_{\mathbb{N}_{0}}$, $\Gamma_{\boldsymbol{m}}$, $\mathbb{W}_{\ell}$, $\chi_{\boldsymbol{m}}$, $\psi_{0}$, $\psi^{m}_{\boldsymbol{m}}$, $J_{\nu}$, $\nu_{0}$, $\|\cdot\|_{t}$ §2.3
$u^{i}$, $\Gamma^{c}$, $P$, $\mathcal{S}$, $S$, $\phi$, $a(\cdot,\cdot)$, $a^{\Omega}(\cdot,\cdot)$, $S^{\Omega}$, $u^{\infty}$, $\Phi^{\infty}$, $\mathds{S}^{n}$ §4
$t_{d}$, $\mathbb{S}$ §4.1
$\{T_{j}\}_{j=1}^{N}$, $\mathbb{V}_{N}$, $V_{N}$, $\phi_{N}$, $A$, $\vec{c}$, $\vec{b}$, $J$ §5
$\mathbb{X}_{\nu}$, $K_{\nu}$, $\mathbb{P}_{\nu}$, $X_{\nu}$, $\mathbb{Y}_{h}$, $Y_{h}$, $L_{h}$, $h$ §5.1
$c_{t}$ §5.3
$A^{Q}$, $\vec{b}^{Q}$, $\vec{c}^{Q}$, $\phi_{N}^{Q}$, $J^{Q}$, $\mathrm{Hull}$ §5.4
Acknowledgements. We are grateful to Martin Costabel for instructive discussions related to this project. SC-W was supported by Engineering and Physical Sciences Research Council (EPSRC) Grant EP/V007866/1, and DH and AG by EPSRC grants EP/S01375X/1 and EP/V053868/1. AM was supported by the PRIN project “NA-FROM-PDEs” and by the MIUR through the “Dipartimenti di Excellenza” Program (2018–2022)–Dept. of Mathematics, University of Pavia. AC was supported by CIDMA (Center for Research and Development in Mathematics and Applications) and FCT (Foundation for Science and Technology) within project UIDB/04106/2020. The authors acknowledge the use of the UCL Myriad High Performance Computing Facility (Myriad@UCL), and associated support services, in the completion of this work.